Purdue University MA375 Discrete Mathematics Spring 2012 Syllabus Instructor: Uli Walther Office: 746 MATH Phone: 49-41959

                                E-mail: walther@math.purdue.edu

1 General Information

1.1 Class time / location:

TTh 10:30-11:45pm, LWSN B151.

1.2 Office Hours:

Tue 12:00-1:00 and Th 1:00-2:00.

1.3 Textbook:

Discrete Mathematics and its Applications Kenneth H. Rosen 7th Edition.)

We will cover material of chapters , 4, 5, 6, 7, 8, 10 and 12.

1.4 Webpage:

Under http://www.math.purdue.edu/~walther/teach, you find this syllabus

as well as some other info concerning the course, including a link to our Project Rhea (see below).

1.5 Statement for Students with Disabilities

See the webpage, http://www.math.purdue.edu/~walther/teach.

1.6 Calculators

You may use calculators as you wish for homework or the tests. On the other hand, numerical answers alone will get you no credit. This class emphasizes understanding the process that leads to a solution, not the solution itself. Credit will be given for the explanation of the method (if correct) and how it is carried out.

2 Course Structure

If you disagree with the grading in any category, you need to resubmit the item in question with a written explanation stating why (and not just that) you deserve more credit.

2.1 Homework:

Homework will be collected every Thursday in class or in my office no later than 2:00pm. No late assignments will be accepted. Neither will homework deposited anywhere else.

The homework due in any given week is the homework corresponding to the material of the previous week. Homework must be readable and must be stapled. Illegible scribblings will receive no credit from the grader.

You are encouraged to attempt all the questions and discuss with your classmates. However, the write-up must be of your own.

• WORK MUST BE SHOWN TO EXPLAIN YOUR ANSWERS. NO CREDIT WILL BE GIVEN FOR JUST THE NUMERICAL AN-SWERS. THIS ALSO APPLIES TO EXAMS.

2.2 Project Rhea

You may earn up to 5% extra credit from your contributions to Rhea. See

 https://www.projectrhea.org/rhea/index.php/2012_Spring_MA_375_Walther

2.3 Midterms:

There will be one midterm, Thursday, March 8. Makeup tests may be given in extraordinary instances, but only with documented reasons.

2.4 Final Exam:

The Final Exam will be given during the Final Exam Week.

The location and date of the final will be announced in class.

2.5 Course Grade:

Your course grade will be determined using the following distributions:

Rhea  5%
HW  30%
Midterm 30%
Final  40%